THE SEARCH FOR CHARGE TRANSFER STATES: BENZONITRILE AND RELATED MOLECULES

\underline{DAVID R. BORST} and DAVID W. PRATT.^{a}, Department of Chemistry, University of Pittsburgh, Pittsburgh, Pennsylvania 15260.

The high resolution electronic spectra of benzonitrile and related molecules have been recorded, including several higher vibronic bands in the S_{1} ← S_{0} system. Analyses of the spectra provide information about the geometries of the different molecular assemblies in both their ground and first excited states. There is no evidence for a charge transfer state up to +703 cm^{−1} above the S_{1} origin transition. Related molecular assemblies will also be discussed.

aWork supported by NSF